Two charges lie on positive x-axis . Charge A (3.0 x 10-10 C) is 3.0 cm from the origin and Charge B is 5.0 cm from the origin (-4.0 x 10-10 C). What is the total force exerted by these two charges on charge C (6.0 x 10-10 C ) located at the origin ?



1
Expert's answer
2022-04-01T16:32:27-0400

Force


"F=\frac{kq_1q_2}{r^2}=\frac{9\times10^{9}\times3\times10^{-10}\times6\times10^{-10}}{0.03^2}=1.8\times10^{-06}N"


"F=\frac{kq_1q_2}{r^2}=-\frac{9\times10^{9}\times4\times10^{-10}6\times10^{-10}}{0.05^2}=-8.64\times10^{-7}N"

Net force

"F=F_1+F_2"

"F=(1.8-0.864)\times10^{-6}N"

"F=9.36\times10^{-7}N"